OpenStudy (lgbasallote):

LGBARIDDLE What do you get when you add 3 to 303 six times?

OpenStudy (zzr0ck3r):

306

OpenStudy (anonymous):

303333333

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

simplest thinker wins

OpenStudy (anonymous):

323036

OpenStudy (anonymous):

321

OpenStudy (anonymous):

303+3*6=321

OpenStudy (zzr0ck3r):

321

OpenStudy (anonymous):

3033333336

OpenStudy (compassionate):

306 + 6 + 6 + 6 + 6 + 6 + 6

OpenStudy (lgbasallote):

seems math lovers really are complicated thinkers

OpenStudy (compassionate):

Wait, I know! You get a number!

OpenStudy (lgbasallote):

not that simple

OpenStudy (compassionate):

Ok. You get lots of numbers.

OpenStudy (zzr0ck3r):

303*6+3

OpenStudy (zzr0ck3r):

1821

OpenStudy (laronjaxon):

is it like 303x6?3x6? or s it 321

hartnn (hartnn):

u get an answer.

OpenStudy (lgbasallote):

here comes @hartnn ....

OpenStudy (lgbasallote):

i thought you were able to solve it...

OpenStudy (compassionate):

The answer is 42!

OpenStudy (lgbasallote):

you're on the right track though @hartnn ...simple-thinking...

OpenStudy (anonymous):

3x6=18 303+18=321

OpenStudy (lgbasallote):

still no

OpenStudy (lgbasallote):

think like a math hater

ganeshie8 (ganeshie8):

you get use to additions

OpenStudy (zzr0ck3r):

i still say it does not matter how many times you add 3 to 303 you always get 306:)

OpenStudy (anonymous):

303+3 then multiply to 6

OpenStudy (compassionate):

I bet we can use the distributive property.

OpenStudy (lgbasallote):

i think @zzr0ck3r is on the right track....

OpenStudy (compassionate):

3 to 303 six times? 3(300 + 3)

OpenStudy (zzr0ck3r):

306*6

OpenStudy (lgbasallote):

hah nope

OpenStudy (lgbasallote):

you were so close

OpenStudy (anonymous):

3(303+3)?

OpenStudy (compassionate):

You will get 603 :D

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

read the question carefully

OpenStudy (zzr0ck3r):

303+3^6

OpenStudy (anonymous):

323*036

OpenStudy (lgbasallote):

exponents starting to come and play too huh

OpenStudy (anonymous):

1836

OpenStudy (anonymous):

1821

OpenStudy (anonymous):

303303303303303303+3?

OpenStudy (lgbasallote):

nope

OpenStudy (compassionate):

9.11.01

OpenStudy (lgbasallote):

that's scary @Compassionate

OpenStudy (anonymous):

6.(303+3)

OpenStudy (anonymous):

32303*6

OpenStudy (lgbasallote):

no...

OpenStudy (anonymous):

306!!!

OpenStudy (anonymous):

you always get it!

OpenStudy (lgbasallote):

this is fun...it's so simple and yet it's so hard

OpenStudy (anonymous):

(303^6)+3 ?

OpenStudy (compassionate):

Maybe he wants us to put it in an integral?

OpenStudy (lgbasallote):

aren't you reppeating what @zzr0ck3r said @RaphaelFilgueiras ?

OpenStudy (anonymous):

303+3=306 303+3=306 ...

OpenStudy (anonymous):

773848189788132

OpenStudy (zzr0ck3r):

thats the first answer of the post, I dont think its what hes looking for

OpenStudy (compassionate):

\[\lim_{303 + 6*3 \rightarrow x}\]

OpenStudy (lgbasallote):

now how did that number came to be

OpenStudy (zzr0ck3r):

now that just does not make sense

OpenStudy (compassionate):

\[\sqrt{303} + \frac{ 303 }{ 2 }(3*6)\]

OpenStudy (zzr0ck3r):

303*6+3 did i do this one yet?

OpenStudy (lgbasallote):

lol like i said simplest thinker wins

OpenStudy (lgbasallote):

that's still not as right as the first ten times it was suggested @zzr0ck3r

OpenStudy (zzr0ck3r):

in other words we are not going to like the abswer:)

OpenStudy (lgbasallote):

probably...

OpenStudy (zzr0ck3r):

hehe

OpenStudy (anonymous):

lol

OpenStudy (compassionate):

You get math

OpenStudy (lgbasallote):

do you like how this looks? 306, 306, 306, 306, 306, 306

OpenStudy (zzr0ck3r):

hahaha:P

OpenStudy (zzr0ck3r):

yes

OpenStudy (lgbasallote):

told you simplest thinker wins

OpenStudy (zzr0ck3r):

I was close

OpenStudy (lgbasallote):

yes...

OpenStudy (zzr0ck3r):

lol